OpenStudy (anonymous):

When f(x) = x^3 - 4x^2 + ax + b is divided by (x+3), the remainder is -64. (x-5) is a factor of f(x). Find the values of a and b. HEEELP PLEASE!

OpenStudy (anonymous):

\[ \frac {x^3 - 4x^2 + ax + b}{ x+3}=21 + a - 7 x + x^2+ \frac {-63 - 3 a + b}{ x+3}\\ -64 = -63 - 3 a + b\\ 3a -b =1 \] Try to find another relation between a and b

OpenStudy (anonymous):

\[ f(5)=0\\ 25 + 5 a + b =0\\ 5 a + b = -25 \]

OpenStudy (anonymous):

Solve for a and b and you find a = -3, b = -10

OpenStudy (alexwee123):

wait :o

OpenStudy (alexwee123):

nm ....

OpenStudy (anonymous):

Your polynomial is \[ f(x) = x^3-4 x^2-3 x-10 \]

OpenStudy (anonymous):

OHHHH ok I see :D thank you so much! I have trouble with long division polynomials -_-

OpenStudy (anonymous):

yw